### v1

# Learn To be Efficient: Build Structured Sparsity in Large Language Models

Abstract
Large Language Models (LLMs) have achieved remarkable success with their billion-level parameters, yet they incur high inference overheads. The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference. Existing methods only focus on utilizing this naturally formed activation sparsity, overlooking the potential for further amplifying this inherent sparsity. In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity. To achieve this, we introduce a novel algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance. Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like GPT and LLaMA with soft activation functions. We evaluate LTE on four models and eleven datasets. The experiments show that LTE achieves a better trade-off between sparsity and task performance. For instance, LTE with LLaMA provides a 1.83x–2.59x FLOPs speed-up on language generation tasks, outperforming the state-of-the-art methods.
1 Introduction
Large language models (LLMs), like GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022a), and LLaMA (Touvron et al., 2023a, b), demonstrated emergent natural language ability as model parameters scale to the billion level. Those works have pushed state-of-the-art performance in a wide range of natural language processing applications. However, as the number of model parameters increases, LLMs also become quite expensive for deployment, challenged by the significant computational and memory costs during inference (Liu et al., 2023). Such expensive costs not only significantly inflate the deployment cost but also affect the user experience in many latency-sensitive applications, such as chatbots (Achiam et al., 2023) and autonomous driving (Li et al., 2023a). Various strategies have been introduced to improve LLM inference efficiency, like quantization (Frantar et al., 2022), model pruning (Ma et al., 2023), token sparsity (Zhang et al., 2023; Xiao et al., 2023), and activation sparsity (Liu et al., 2023; Zhang et al., 2022b).
During LLM inference, Feed Forward Network (FFN) layers have been one primary bottleneck, taking over the FLOPs and IOs (Liu et al., 2023). Recent works (Li et al., 2023b; Liu et al., 2023) report that ReLU-based LLMs have extremely high activation sparsity in those FFN layers. For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023). In light of this favorable balance between accuracy and speedup, existing works explore grouping intermediate neurons in FFN layers into structured chunks (Mirzadeh et al., 2023; Liu et al., 2023), thus achieving hardware-friendly acceleration on modern hardware like GPU (Liu et al., 2023).
Previous efforts focus on utilizing this activation sparsity of ReLU-based LLMs by MoEfication (Zhang et al., 2022b). Given an activation-sparse model, state-of-the-art (SOTA) methods (Zhang et al., 2022b; Liu et al., 2023) propose to convert FFN layers to MoE layers by grouping neurons in the FFN intermediate layer into experts and then applying a router to select most important experts for each input token. Unlike training the MoE model from scratch, which requires a huge amount of resources for pretraining, MoEfication converts FFN layers in pretrained LLMs into MoE layers with relatively small training resources. However, SOTA methods only focus on utilizing this naturally formed activation sparsity of LLMs, overlooking the potential for further amplifying/increasing this inherent sparsity. Moreover, for LLMs with soft activation functions, such as SwiGLU (Shazeer, 2020) in LLaMA, existing works have to replace activation with ReLU to enable sparsity. Unfortunately, our evaluations across many datasets show that this replacement can hurt model performance. Consequently, we need a novel method to improve LLM sparsity under soft activation functions.
In this paper, we hypothesize that LLMs can learn to be efficient and achieve more structured activation sparsity. As shown in Figure 1, our key insight is that, besides optimizing models for task performance alone, we can also train LLMs to be efficiency-aware by developing more structured sparsity, which can be utilized more effectively. However, creating structured sparsity in pretrained LLMs is a non-trivial task due to three practical challenges:
1.
How to group neurons into experts? Since neurons in the same expert are always activated together, a smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
2.
How to train routers more stably? The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2). How to jointly train routers and model in a MoEfication setting is still under-explored.
3.
How to select the right experts in serving? The number of experts needed in MoE layers depends on specific inputs and layers, implying a trade-off between model efficiency and quality.
To address these challenges, we introduce Learn-To-be-Efficient (LTE), a novel training algorithm to train efficiency-aware models. LTE integrates an efficiency loss penalty, encouraging models to activate fewer neurons in their FFN layers. Additionally, LTE adopts a threshold-based Sigmoid routing strategy to select experts and employs a two-stage training mechanism to improve the training stability, achieving a more flexible selection of experts instead of selecting a fixed number of experts for all layers and inputs.
To evaluate the effectiveness of our method, we implement LTE on both encoder-based models (RoBERTa, RoBERTa (Liu et al., 2019)) and decoder-based models (GPT2-Medium (Radford et al., 2019), LLaMA-7B (Touvron et al., 2023a)) with the HuggingFace’s transformers, and evaluate our method on eleven datasets from both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks. For NLU tasks, our experiments show that LTE achieves 80% - 95% sparsity without performance drops. Similarly, LTE also achieves a better trade-off between sparsity and task performance than other baselines. For instance, LLaMA-7B with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks, outperforming all other baseline methods. Our evaluation results demonstrate that LTE effectively builds more structured sparsity even on LLMs with soft activation functions.
2 Related Work
Mixture of Experts (MoE). MoE was proposed by Jacobs et al. a few decades ago to build an integral system with subset networks. Recently, MoE layers have been used in the Transformer architecture as a substitute for the MLP block (Shazeer et al., 2018; Fedus et al., 2022), with the recent Mistral-7B model (Jiang et al., 2023) as a prominent example. In MLP blocks of MoE transformers, instead of using an FFN layer for calculating output, MoE layers construct multiple FFN layers and employ a router to choose a subset of experts to do conditional computation. Even though transformers can have billions of parameters, only a subset of experts are activated for each token, thus effectively increasing model activation sparsity (i.e., the number of skipped neurons in execution).
LLM Sparsity. Recent studies (Li et al., 2023b; Liu et al., 2023) show that trained ReLU-based transformers naturally show a great sparsity in the activation of FFN layers. For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023). While SOTA models are increasingly using diverse soft activations, like GeLU, SiLU, and SwiGLU, existing work show that replacing those soft activations with ReLU and fine-tuning the model can significantly increase the activation sparsity at the cost of a slight performance drop (Zhang et al., 2022b; Mirzadeh et al., 2023; Li et al., 2023b). However, our evaluation shows that this benefit does not always hold for all datasets (Section 5.2).
Speedup from Sparsity. The emerging sparsity indicates that a large portion of neurons are unnecessary in LLMs inference, which can be taken advantage to improve LLM inference efficiency. One way to utilize sparsity is MoEfication (Zhang et al., 2022b). Recent papers (Liu et al., 2023; Zhang et al., 2022b) show that a single FFN layer in pretrained LLMs can be converted to an MoE layer by splitting the matrices of FFN layers into exclusive sub-matrices and employs a router to activate only neurons in a subset of experts. Similar to MoE models, MoEfication can effectively accelerate LLM inference by only using part of the parameters (Zhang et al., 2022b; Liu et al., 2023).
3 Background and Motivation
In this section, we first briefly introduce the MoEfication of FFN layers in transformers. Then, we present a study on the limitation of applying noisy top-K Softmax routing to train transformers for better sparsity.
3.1 Background: MoEfication
FFN layers occupy 60% of the total model computation (Liu et al., 2023) but exhibit high activation sparsity. In contrast to unstructured sparsity methods like weight pruning, activation sparsity in ReLU-based FFN layers is more hardware-friendly, since intermediate neurons in FFN layers can be grouped as structured chunks (Mirzadeh et al., 2023; Liu et al., 2023). MoEfication (Zhang et al., 2022b) is a way to group functional partitions in FFN layers, thereby converting FFN layers to MoE layers for better utilizing the activation sparsity. As known, if a transformer’s hidden states and FFN intermediate dimension are and , respectively, the FF layers process the input as following:
where , are weight metrics, are biased term and is an activation function111SwiGLU in LLaMA needs two matrices to calculate , and the neurons in FFNs can still be used to form experts.. MoEfication aims to group neurons in the FFN intermediate layer into experts. Then, a router is trained for each FFN layer to select most important experts to improve inference sparsity. A recent work, Deja Vu (Liu et al., 2023), also uses a similar setting but treats each neuron as an expert.
However, SOTA MoEfication methods overlook the potential for further optimizing the activation sparsity of LLMs. In this paper, besides optimizing LLMs for task performance, we propose to optimize models for fast inference as well (Figure 1). By designing efficiency-related loss, we find that our novel training method drives LLMs to achieve better sparsity with marginal performance drops. Besides, for LLMs employing soft activations such as GeLU (Hendrycks & Gimpel, 2016) and SwiGLU (Shazeer, 2020), MoEfication (Zhang et al., 2022b) proposes to replace soft activations with ReLU and fine-tune the model to improve activation sparsity. Although recent works (Zhang et al., 2022b; Mirzadeh et al., 2023) show that replacing soft activation functions with ReLU only has a marginal impact on model performance, our evaluations indicate that this conclusion is not universally applicable across all datasets. It also is crucial to develop methods for effectively utilizing sparsity in LLMs with soft activation functions.
3.2 Limitations of Noisy Top-K Softmax Routing
A potential choice to train the model to adapt to MoE layers is noisy top-K Softmax routing (Fedus et al., 2022; Shazeer et al., 2016), which is a commonly adopted approach to train MoE models from scratch, like Switch Transformers (Fedus et al., 2022). This method entails selecting experts based on the highest router outputs and then calculating Softmax values for these selected outputs. The output of the MoE layer is determined by summing the products of these softmax values with their corresponding experts’ outputs, which is critical for differentiating router parameters. Additionally, a small Gaussian noise is added to the router’s outputs during training to encourage exploration across different experts, which addresses the issue of unselected experts being non-differentiable. However, our empirical findings surprisingly indicate that noisy top-K does not converge well and encounters a biased expert score issue in MoEfication tasks.
To study the effectiveness of noisy top-K Softmax routing, we follow the standard MoEfication setting (Zhang et al., 2022b) to form experts in FFN layers: we fine-tune a model on a downstream dataset and group neurons into experts via parameter clustering. Subsequently, we integrate Softmax routers into models and jointly train both the model and routers with noisy top-K routing. The evaluation results on two datasets are shown in Figure 2. Surprisingly, we observe similar accuracy drops across all sparsity levels, including at very low levels of sparsity. To study the reason behind this, we analyze the expert scores produced by routers and uncover a biased expert issue in MoE layers: one expert in each MoE layer receives an expert score close to 1, while the expert scores for all other routers are nearly 0. This means that only one expert contributes to the inference in each MoE layer, which causes the performance drops.
Before MoEfication, expert outputs in FFN layers can be considered to be multiplied by an expert score equal to 1. The sum of expert scores in Softmax routing is constrained to 1, which results in many experts’ outputs being multiplied by expert scores significantly smaller than 1. Meanwhile, fine-tuning LLMs involves a relatively low learning rate and fewer training iterations. This makes it difficult for neurons in FFN layers to adapt to such largely reduced expert scores. Consequently, the model tends to rapidly converge towards allocating the entire Softmax budget to the most important expert, assigning other experts with expert scores close to 0.
Another limitation of the noisy top-K routing approach is its requirement for a fixed, pre-defined number of selected experts for all inputs and layers. However, the sparsity level can vary depending on the specific inputs and layers. A more adaptive expert selection strategy helps achieve a better trade-off between model efficiency and quality.
4 Methodology: Learn To be Efficient
In this section, we present, LTE (Learn To be Efficient), to learn more structured MoE in FFN layers for fast inference.
Challenges. Converting a pre-trained LLM’s MLP block into a structured Mixture of Experts (MoE) leads to three practical challenges:
1.
How to group neurons into experts? Since neurons in the same expert are always activated together, A smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
2.
How to train routers more stably? The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2). How to jointly train routers and model in a MoEfication setting is still under-explored.
3.
How to select the right experts in serving? The number of experts needed can depend on specific inputs and layers. A more adaptive selection strategy leads to a better trade-off between model efficiency and quality.
4.1 Experts Grouping
Expert grouping aims to group neurons that are often activated together, thereby improving efficiency by reducing the total number of activated neurons. Here, constructing too many experts (e.g., each neuron as an expert) can significantly increase the cost, while too few can lead to poor model quality. Moreover, we need to group neurons without a significant performance drop on the pre-trained model.
Following the previous work (Zhang et al., 2022b), we use parameter clustering to split MLP neurons into different experts. Fundamentally, in FFN layers, each neuron is associated with a column of (which is a vector). Parameter clustering split first treats the corresponding vector in as the feature for a neuron, then applies the balanced K-Means algorithm (Malinen & Fränti, 2014) to cluster neurons into clusters. Each cluster has neurons by default and will be treated as an individual MoE expert.222 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: We use the weight of the gate network, , to cluster neurons.
4.2 Adaptive Expert Routing
After allocating neurons to different experts, we need to train and orchestrate routers to select experts for input tokens.
4.2.1 Router structure
Similar to existing MoE designs, we use a single fully-connected layer as the router network. The input to the router is the output of the preceding self-attention block, and the router output is the expert score for each expert.
Routing function. To address the biased expert score issue (Section 3.2), instead of using conventional Softmax routing, we employ the Sigmoid function on router outputs to get expert scores:
where is the input to the corresponding FFN layer, is the index for the expert, and is the router network weight for the -th expert. Unlike the Softmax function, the Sigmoid function computes each expert score independently, which circumvents the biased expert score due to the constraint on the sum of outputs in the Softmax function.
Threshold-based experts selection. SOTA methods (Zhang et al., 2022b; Fedus et al., 2022) select a fixed number of experts for all layers and inputs. However, the necessary number of experts may differ depending on inputs and layers. Here, we propose to select experts more adaptively for different inputs and layers for a better trade-off between model sparsity and quality. Specifically, we leverage a threshold-based method to enable adaptive selection:
where is the output for the -th expert, is the indicator function, and is a predefined threshold. Only the experts whose score is larger than will be activated. Threshold-based routing does not have to select a fixed number of experts in each MoE layer. Instead, it selects experts with sufficiently large expert scores. Consequently, within the same MoE layer, as the router generates larger expert scores, more experts are selected.
4.2.2 Two-stage LTE Training
In practice, training the threshold-based Sigmoid router leads to three challenges. Firstly, Sigmoid function considers individual experts independently, making it hard to tease apart important experts from the rest. Secondly, the threshold-based indicator function in Equation 2 is non-differentiable, thereby preventing the training on routers. Lastly, manually choosing a threshold for each MoE layer can be impractical.
To tackle these challenges, we next propose a novel two-stage training algorithm with an efficiency loss penalty to MoEfy LLMs. As shown in Algorithm 1, our training algorithm consists of two training stages: a) model-router training stage and b) model adaption stage.
Stage 1: Model-router training. The goal of the model-router joint training stage is to train Sigmoid routers to capture the importance of experts for given inputs. In this stage, we jointly train routers and model parameters.
To address the non-differentiability issue, we switch the expert selection to a “soft” mode in this stage:
Instead of selecting a subset of experts, we always select all experts and multiply expert outputs with the corresponding expert score to make both router and model differentiable. Since there is no discrete selection in MoE layers, all parameters are trained for each input token.
Compared to Softmax routers, Sigmoid routers score each expert independently and do not introduce any competition on expert scores among experts. This makes Sigmoid routers alone cannot identify more important experts in MoE layers. As such, we design an efficiency loss penalty, , to introduce competition among experts for Sigmoid routers:
The efficiency loss calculates the mean of expert scores across all layers. This efficiency loss penalizes the output magnitudes of routers, driving routers to selectively allocate lower expert scores to less important experts, which helps distinguish experts with different importance. Moreover, instead of assigning the same sparsity for all layers, the efficiency loss also induces inter-layer competition, allowing LTE models to allocate adaptive sparsity for different layers. Our analysis in Figure 6 shows that LTE models are capable of adaptively allocating sparsity across different layers.
For the threshold in Equation 2, instead of choosing a threshold for each router, we propose to select a predetermined fixed threshold and then train models to fit this threshold, i.e., train models to be threshold-aware. Specifically, we use a threshold separability regularizer to drive models to make expert scores more separable for this given threshold:
where is the number of layers, and is a predefined threshold for all routers (we set for all experiments). The separability loss encourages router outputs to diverge from the threshold , which makes the router outputs more separable. An example is illustrated in Figure 3. The only difference between the models in the two figures is if the model is trained with the separability loss. We find that, when training without the separability loss, the expert scores are hard to separate, making it hard to decide the threshold. However, training with the separability loss markedly improves the separability of expert scores, allowing us to choose experts with the predefined threshold.
Combined with the task performance loss , our training loss in model-router training state is:
The efficiency coefficient, , is used to control the trade-off between inference efficiency and task performance. A higher indicates a larger inference cost penalty, leading to a more sparse model.
Different from the pretrained LLMs, MoE routers are randomly initiated. We use a larger learning rate (0.01) for routers’ parameters to make routers converge faster.
Stage 2: Model Adaptation. To accelerate model inference, we need to switch routers to discrete selection mode (Equation 2), and the model trained in Stage 1 needs further training to adapt to the changes in router outputs. In the model adaptation stage, we freeze the router parameters, switch routers to discrete selection mode (Equation 2), and fine-tune the model to adapt to the discrete routing with the task performance loss: .
5 Experiment
In this section, we evaluate the effectiveness of LTE with different models on various datasets. In Section 5.2, the experiment results on NLU tasks show that LTE models achieve 80 - 95% sparsity in FFN layers without performance drop. Similarly, LTE also outperforms all other baselines on NLG tasks. For instance, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks.
5.1 Experimental Setting
Models. We implement LTE on both encoder-based models (RoBERTa, RoBERTa) and decoder-based models (GPT2-Medium, and LLaMA-7B) with the HuggingFace’s transformers. Following previous work (Zhang et al., 2022b), we set each expert to contain 32 neurons in FFN layers for all models.
Datasets. We evaluate LTE performance using 11 datasets across two setups:
Natural Language Understanding (NLU): we evaluate on eight tasks from the GLUE dataset (Wang et al., 2018) including sentiment classification:SST-2 (Socher et al., 2013), entailment: RTE (Dagan et al., 2005), grammatical classification: CoLA (Warstadt et al., 2018) two natural language reference tasks: MNLI (Williams et al., 2018), QNLI (Rajpurkar et al., 2016), and three tasks measuring sentence similarity: QQP (Iyer et al., 2017), STS-B (Cer et al., 2017), MPRC (Dolan & Brockett, 2005). We report the Pearson correlation coefficient for STS-B, the Matthews correlation coefficient for CoLA, and the accuracy for the other six datasets.
Natural Language Generation (NLG): we evaluate on data-to-text (E2E Challenge (Novikova et al., 2017)), text-to-text summarization (XSum (Narayan et al., 2018)), and next token prediction language modeling (Wikitext103 (Merity et al., 2016)). For both E2E and XSum, we report the ROUGE-L (Lin, 2004) to measure the generation text quality and report the perplexity performance for the Wikitext dataset.
Baselines. (1) MoEfication (Zhang et al., 2022b; Liu et al., 2023): MoEfication proposes to replace soft activations with ReLU and then fine-tune the model on downstream datasets. In our evaluation, we report the number of MoEfication with parameter clustering and ground-truth expert selection. Note that ground-truth selection is not practically feasible but can be considered as an upper bound for expert selection. (2) Keep-the-Largest-Activation (KLA): To explore the sparsity upper-bound in LLMs trained with task performance loss alone, we manually create an upper-bound baseline based on an ideal scenario. We assume that an oracle can always predict the activation value of each neuron with no cost. Given a sparsity level, the oracle always selects neurons with the largest absolute activation values. Although KLA is not feasible in real-world settings, it represents an upper-bound performance for models that do not optimize efficiency in training. The models used in KLA are also models fine-tuned on downstream tasks.
5.2 Performance Comparison
No performance drop with 80-95% sparsity on NLU tasks. We first evaluate our methods on eight NLU Tasks in GLUE datasets(Wang et al., 2018) with two encoder-based models (RoBERTa and RoBERTa). Due to the space limit, we report four datasets (MRPC, SST2, QNLI, and MNLI) in Figure 4 and other four datasets (RTE, CoLA, STSB, and QQP) in Figure 8 in Appendix B.
As shown in Figure 4, our method (LTE) consistently outperforms the baseline models on all datasets. Both MoEfication and KLA exhibit performance drop when sparsity reaches 70%–80%, while LTE maintains good performance even at higher sparsity levels (90%). The major reason behind this performance gap is that, although LLMs naturally have high sparsity, they are still not optimized for efficiency. The efficiency loss penalty in LTE training loss effectively enables LLMs to activate fewer neurons while minimally impacting task performance.
Moreover, we notice that while ReLU-based models achieve high sparsity with a slight decrease in performance, replacing GeLU with ReLU in RoBERTa can affect performance on some datasets. For instance, replacing GeLU with ReLU in a RoBERTa model results in a 10% accuracy drop on the MRPC dataset.
50% FLOPs saving on NLG tasks. We next evaluate LTE with decoder-based models on three types of generation tasks. Similar to NLU tasks, Figure 5 shows that our method outperforms all other baselines. NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level compared to NLU tasks. However, LTE-trained models have better resilience to the sparsity increase. Especially at a high sparsity level, other baselines fail to generate meaningful content, but LTE still generates content with a marginal quality drop. In NLG tasks, we also observe that replacing soft activation with ReLU can hurt model performance. However, training with ReLU-LLaMA largely mitigates some of the performance drops.
| XSum | E2E | WikiText-103 | |
| Full | 12.06 | 12.06 | 12.06 |
| KLA | 7.14 | 6.42 | 7.87 |
| MoEfication | 10.45 | N/A | N/A |
| R-LLaMA+MoE | 8.27 | 7.39 | 11.1 |
| LTE (Ours) | 5.38 | 4.65 | 6.59 |
To better understand the computation saving of LTE, we compare the FLOPs per token of different methods on LLaMA, allowing an up to 0.05 ROUGE-L decrease for XSum and E2E, and an up to 0.5 perplexity (PPL) increase for WikiText-103. It is noteworthy that the FLOPs for routers are included in our method, constituting approximately 1% of the total FLOPs of FFN layers. The results are presented in Table 1. The evaluation results show that LTE provides a 1.83x - 2.59x FLOPs speed-up and gives the most FLOPs savings among all methods. Those evaluation results demonstrate that LTE effectively provides higher sparsity for LLMs even with soft activation functions.
5.3 Ablation Study & Analysis
Adaptive Sparsity across Layers. Instead of assigning a fixed sparsity to each layer, LTE uses the efficiency loss to introduce competition across different layers to allocate the computation budget to more important layers. Here, we conduct a study on the distribution of sparsity across layers in models trained with LTE. As shown in Figure 6, we observe that LTE achieves different sparsity levels across different layers. For example, the MRPC (RoBERTa) model with 0.46 sparsity has 0.2 sparsity at layer 2, but has more than 0.8 sparsity at layers 10 and 11.
Another intriguing finding is that RoBERTa and GPT2-M have very different sparsity patterns. In RoBERTa, the sparsity of MoE layers significantly increases with depth, contrasting with GPT2-M, where MoE layer sparsity decreases as layer depth increases. This difference can stem from differences in the design of these two transformers and tasks. Given that MRPC is a sentence classification task, encoder-based transformers like RoBERTa primarily rely on the ”[CLS]” token for classification. Consequently, other tokens become less important in deeper layers, allowing sparser FFN layers. In contrast, GPT2, a decoder-based transformer designed for language generation, requires all tokens to generate the next token. Furthermore, recent research (Geva et al., 2021) on interpretability suggests that deeper FFN layers store more complex semantic information, which may drive deep FFN layers to be less sparse.
Effectiveness of LTE-trained routers. We next demonstrate the effectiveness of LTE’s first training stage, the model-router training stage, in helping LLMs gain better sparsity. To build a baseline, instead of using routers trained in the model-router training stage, we use randomly initialized networks as routers and train LLMs to adapt to random routers. Since we cannot control the sparsity of LLMs with a threshold in a random router, we decide the sparsity for MoE layers by picking K top experts: we first fix a K to decide the sparsity and then start to train the model. Figure 7 compares the performance of models with different routers. We observe that routers trained with LTE significantly contribute to achieving a better trade-off between sparsity and model performance, which proves the effectiveness of the model-router training stage.
6 Conclusion
In this paper, we explore how to enable better activation sparsity for fast LLM inference. We introduce a novel algorithm, LTE, for the model to automatically learn to activate fewer neurons, by grouping neurons into different experts and then adaptively selecting important ones for specific inputs and layers. Our extensive evaluations, spanning four LLMs and eleven datasets, show that LTE can achieve 1.83x - 2.59x FLOPs speed-up on LLaMA, thus execution efficiency, outperforming state-of-the-art designs. We believe that our work can inspire more research on designing more efficiency-aware training methods, making LLMs more accessible to the broad community.
Impact Statement
This paper presents the Learn-To-be-Efficient (LTE) algorithm to accelerate the LLM inference. A more efficient LLM inference can significantly enhance the accessibility and sustainability of LLMs. By reducing computational demands, it addresses environmental concerns through lower energy consumption and helps democratize access to advanced AI technologies. We believe that our work can better help smaller organizations, educational institutions, and researchers, who previously faced barriers due to resource limitations, access LLMs more easily. To sum up, the LTE algorithm not only advances the field technically but also broadens the scope of AI’s benefits to a wider, more inclusive community.
Appendix A Experiment Settings
We presented all training hyperparameters in Table 2, 3, 4, and 5. For hyperparameters presented in {}, we select the best hyperparameter for each task. We follow the settings in RoBERTa paper (Liu et al., 2019) to fine-tune RoBERTa on GLUE datasets. We set the coefficient for the separability loss ( in Equation 6) to be 0.5 for all stage 1 training.
Appendix B Additional Evaluation Results
In Figure 8, we present the evaluation results of the rest four NLU datasets. We have similar findings as we discussed in Section 5.2: LTE achieves a better trade-off between sparsity and task performance. However, we notice that, LTE outperforms MoEfication on the CoLA dataset, but has worse performance than KLA at low sparsity level. The size of CoLA is relatively small and may not be efficient for models to learn good routers.
| Hyperparameter | LTE-Stage 1 | LTE-Stage 2 |
| Learning rate | {2e-5, 5e-5} | {5e-5} |
| Training batch size | {16, 32} | {16, 32} |
| Training epochs | {1, 10} | {1, 3, 10} |
| Weight decay | 0 | 0 |
| Warm up ratio | 0.06 | 0.06 |
| Hyperparameter | LTE-Stage 1 | LTE-Stage 2 |
| Learning rate | {2e-6, 5e-6, 3e-5} | {2e-6, 5e-6, 3e-5} |
| Training batch size | {16, 32} | {16, 32} |
| Training epochs | {1, 2, 3, 10} | {3, 5, 10} |
| Weight decay | 0 | 0 |
| Warm up ratio | 0.06 | 0.06 |
| Hyperparameter | Fine-tuning | LTE-Stage 1 | LTE-Stage 2 |
| Learning rate | 5e-5 | 5e-5 | 5e-5 |
| Training batch size | 8 | 8 | 8 |
| Training epochs | 3 | 1 | (1, 1, 3) |
| Weight decay | 0 | 0 | 0 |
| Warm up ratio | 0.06 | 0.06 | 0.06 |
| Hyperparameter | Fine-tuning | LTE-Stage 1 | LTE-Stage 2 |
| Learning rate | 1.5e-5 | 1.5e-5 | 1.5e-5 |
| Training batch size | 16 | 16 | 16 |
| Training epochs | 3 | 1 | 1 |
| Weight decay | 0 | 0 | 0 |
| Warm up ratio | 0.06 | 0.06 | 0.06 |

### v3

@@
 The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference.
-Existing methods only focus on utilizing this naturally formed activation sparsity, overlooking the potential for further amplifying this inherent sparsity.
+However, existing methods only focus on utilizing this naturally formed activation sparsity in a post-training setting, overlooking the potential for further amplifying this inherent sparsity.
 In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity.
-To achieve this, we introduce a novel algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance.
-Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like GPT and LLaMA with soft activation functions.
-We evaluate LTE on four models and eleven datasets.
-The experiments show that LTE achieves a better trade-off between sparsity and task performance.
-For instance, LTE with LLaMA provides a 1.83x–2.59x FLOPs speed-up on language generation tasks, outperforming the state-of-the-art methods.
+To achieve this, we introduce a novel training algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance.
+Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like LLaMA using non-ReLU activations.
+Extensive evaluation on language understanding, language generation, and instruction tuning tasks show that LTE consistently outperforms SOTA baselines.
+Along with our hardware-aware custom kernel implementation, LTE reduces LLaMA2-7B inference latency by 25% at 50% sparsity.
 1 Introduction
-Large language models (LLMs), like GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022a), and LLaMA (Touvron et al., 2023a, b), demonstrated emergent natural language ability as model parameters scale to the billion level.
-Those works have pushed state-of-the-art performance in a wide range of natural language processing applications.
-However, as the number of model parameters increases, LLMs also become quite expensive for deployment, challenged by the significant computational and memory costs during inference (Liu et al., 2023).
-Such expensive costs not only significantly inflate the deployment cost but also affect the user experience in many latency-sensitive applications, such as chatbots (Achiam et al., 2023) and autonomous driving (Li et al., 2023a).
-Various strategies have been introduced to improve LLM inference efficiency, like quantization (Frantar et al., 2022), model pruning (Ma et al., 2023), token sparsity (Zhang et al., 2023; Xiao et al., 2023), and activation sparsity (Liu et al., 2023; Zhang et al., 2022b).
-During LLM inference, Feed Forward Network (FFN) layers have been one primary bottleneck, taking over the FLOPs and IOs (Liu et al., 2023).
-Recent works (Li et al., 2023b; Liu et al., 2023) report that ReLU-based LLMs have extremely high activation sparsity in those FFN layers.
-For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023).
-In light of this favorable balance between accuracy and speedup, existing works explore grouping intermediate neurons in FFN layers into structured chunks (Mirzadeh et al., 2023; Liu et al., 2023), thus achieving hardware-friendly acceleration on modern hardware like GPU (Liu et al., 2023).
-Previous efforts focus on utilizing this activation sparsity of ReLU-based LLMs by MoEfication (Zhang et al., 2022b).
-Given an activation-sparse model, state-of-the-art (SOTA) methods (Zhang et al., 2022b; Liu et al., 2023) propose to convert FFN layers to MoE layers by grouping neurons in the FFN intermediate layer into experts and then applying a router to select most important experts for each input token.
-Unlike training the MoE model from scratch, which requires a huge amount of resources for pretraining, MoEfication converts FFN layers in pretrained LLMs into MoE layers with relatively small training resources.
-However, SOTA methods only focus on utilizing this naturally formed activation sparsity of LLMs, overlooking the potential for further amplifying/increasing this inherent sparsity.
-Moreover, for LLMs with soft activation functions, such as SwiGLU (Shazeer, 2020) in LLaMA, existing works have to replace activation with ReLU to enable sparsity.
-Unfortunately, our evaluations across many datasets show that this replacement can hurt model performance.
-Consequently, we need a novel method to improve LLM sparsity under soft activation functions.
+Large language models (LLMs), like GPT-3 [2], OPT [42], and LLaMA [35, 36], have demonstrated impressive natural language ability.
+However, the skyrocketing number of model parameters has not only inflated the deployment costs of LLMs, due to their significant computational and memory demands during inference [20], but also affected the user experience in many latency-sensitive applications, such as chatbots [1] and autonomous driving [16].
+Recent advances have been leveraging sparsity to improve LLM inference efficiency, including model weight quantization [8] and pruning [21], token sparsity [44, 41], and activation sparsity [20, 43].
+During LLM inference, Feed Forward Network (FFN) layers are the primary efficiency bottleneck, accounting for over 60% of the FLOPs and I/O operations [20], but they exhibit high activation sparsity, especially in ReLU-based LLMs [17].
+For example, in a 175B OPT model, more than 95% of activations in FFN layers are zeros [20].
+Recent advances leverage this activation sparsity via MoEfication [43, 24]: they convert FFN layers to MoE layers by grouping neurons in the FFN intermediate layer into experts and then applying a router to select most important experts for each input token.
+Unlike training the MoE model from scratch, MoEfication converts FFN layers in pretrained LLMs into MoE layers with relatively small training resources.
+However, existing advances only focus on manipulating the pre-trained activation sparsity of neurons, overlooking the potential for further increasing this inherent sparsity.
+Moreover, they are limited to the historically ReLU-based LLMs, whereas emerging advanced models use soft activation functions for better model quality (e.g., SwiGLU in LLaMA [29] and GeGLU in Gemma [34]), exhibiting much lower natural activation sparsity.
+While existing works suggest replacing the model activation with ReLU to enable sparsity [43, 24], we show that this can hurt model performance (Section 5.2).
 In this paper, we hypothesize that LLMs can learn to be efficient and achieve more structured activation sparsity.
-As shown in Figure 1, our key insight is that, besides optimizing models for task performance alone, we can also train LLMs to be efficiency-aware by developing more structured sparsity, which can be utilized more effectively.
-However, creating structured sparsity in pretrained LLMs is a non-trivial task due to three practical challenges:
+As shown in Figure 1, our key insight is that, without compromising model quality, we can also train LLMs to be efficiency-aware by developing more structured sparsity, which is more friendly for achieving hardware speedup.
+However, creating structured sparsity in pretrained LLMs is non-trivial due to two practical challenges:
 1.
-How to group neurons into experts?
-Since neurons in the same expert are always activated together, a smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
+How to train routers more stably?
+The widely-used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2).
+How to jointly train the model and routers in a MoEfication setting is still open-ended.
 2.
-How to train routers more stably?
-The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2).
-How to jointly train routers and model in a MoEfication setting is still under-explored.
-3.
 How to select the right experts in serving? The number of experts needed in MoE layers depends on specific inputs and layers, implying a trade-off between model efficiency and quality.
 To address these challenges, we introduce Learn-To-be-Efficient (LTE), a novel training algorithm to train efficiency-aware models.
-LTE integrates an efficiency loss penalty, encouraging models to activate fewer neurons in their FFN layers.
-Additionally, LTE adopts a threshold-based Sigmoid routing strategy to select experts and employs a two-stage training mechanism to improve the training stability, achieving a more flexible selection of experts instead of selecting a fixed number of experts for all layers and inputs.
-To evaluate the effectiveness of our method, we implement LTE on both encoder-based models (RoBERTa, RoBERTa (Liu et al., 2019)) and decoder-based models (GPT2-Medium (Radford et al., 2019), LLaMA-7B (Touvron et al., 2023a)) with the HuggingFace’s transformers, and evaluate our method on eleven datasets from both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks.
-For NLU tasks, our experiments show that LTE achieves 80% - 95% sparsity without performance drops.
-Similarly, LTE also achieves a better trade-off between sparsity and task performance than other baselines.
-For instance, LLaMA-7B with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks, outperforming all other baseline methods.
+LTE integrates an efficiency loss penalty, encouraging models to activate fewer neurons in their FFN layers while keeping good task performance.
+Additionally, LTE adopts a threshold-based Sigmoid routing strategy to select experts and employs a two-stage training mechanism to improve training stability.
+LTE achieves a more flexible selection of experts instead of selecting a fixed number of experts for all layers and inputs.
+Same as Deja Vu [20] and moefication [43], LTE provides very structured sparsity.
+We further develop a custom CUDA kernel with Triton, a Python-like CUDA programming language, to enable wall-clock time speedup from such structured sparsity (Section 4.3).
+We evaluate LTE on both encoder-based models (RoBERTa, RoBERTa [19]) and decoder-based models (GPT2-Medium [27], LLaMA2-7B [35]) with the HuggingFace’s transformers.
+Our extensive experiments on Natural Language Understanding (NLU) tasks, downstream Natural Language Generation (NLG) tasks, and instruction tuning tasks, show that LTE consistently outperforms state-of-the-art designs.
+For instance, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks without compromising model quality.
+After integrating our hardware-efficient implementation of sparse matrix multiplication kernel, LTE reduces LLaMA2-7B 25% wall-clock time latency at around 50% sparsity.
 Our evaluation results demonstrate that LTE effectively builds more structured sparsity even on LLMs with soft activation functions.
@@
 Mixture of Experts (MoE).
-MoE was proposed by Jacobs et al.
-a few decades ago to build an integral system with subset networks.
-Recently, MoE layers have been used in the Transformer architecture as a substitute for the MLP block (Shazeer et al., 2018; Fedus et al., 2022), with the recent Mistral-7B model (Jiang et al., 2023) as a prominent example.
-In MLP blocks of MoE transformers, instead of using an FFN layer for calculating output, MoE layers construct multiple FFN layers and employ a router to choose a subset of experts to do conditional computation.
+MoE was proposed by [13] a few decades ago to build an integral system with subset networks.
+Recently, MoE layers have been used in the Transformer architecture as a substitute for the MLP block [30, 6], with the recent Mistral-7B model [14] as a prominent example.
+In MLP blocks of MoE transformers, instead of using an FFN layer for calculating output, MoE layers construct multiple smaller FFN layers and employ a router to choose a subset of experts to do conditional computation.
 Even though transformers can have billions of parameters, only a subset of experts are activated for each token, thus effectively increasing model activation sparsity (i.e., the number of skipped neurons in execution).
-LLM Sparsity.
-Recent studies (Li et al., 2023b; Liu et al., 2023) show that trained ReLU-based transformers naturally show a great sparsity in the activation of FFN layers.
-For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023).
-While SOTA models are increasingly using diverse soft activations, like GeLU, SiLU, and SwiGLU, existing work show that replacing those soft activations with ReLU and fine-tuning the model can significantly increase the activation sparsity at the cost of a slight performance drop (Zhang et al., 2022b; Mirzadeh et al., 2023; Li et al., 2023b).
-However, our evaluation shows that this benefit does not always hold for all datasets (Section 5.2).
-Speedup from Sparsity.
-The emerging sparsity indicates that a large portion of neurons are unnecessary in LLMs inference, which can be taken advantage to improve LLM inference efficiency.
-One way to utilize sparsity is MoEfication (Zhang et al., 2022b).
-Recent papers (Liu et al., 2023; Zhang et al., 2022b) show that a single FFN layer in pretrained LLMs can be converted to an MoE layer by splitting the matrices of FFN layers into exclusive sub-matrices and employs a router to activate only neurons in a subset of experts.
-Similar to MoE models, MoEfication can effectively accelerate LLM inference by only using part of the parameters (Zhang et al., 2022b; Liu et al., 2023).
+LLM Contextual Sparsity.
+Recent studies [17, 20] show that trained ReLU-based transformers naturally show a great sparsity in the activation of FFN layers.
+For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros [20].
+While SOTA models are increasingly using diverse soft activations, like GeLU and SwiGLU, existing work shows that replacing soft activations with ReLU and fine-tuning the model can increase the activation sparsity without greatly hurting model quality [43, 24, 17].
+Yet, we find that this benefit is not consistent for all datasets (Section 5.2).
+This emerging sparsity indicates that a large portion of neurons are unnecessary in LLM inference (i.e., sparsity), which we can leverage to improve LLM inference efficiency like using MoEfication [43].
+Recent papers [20, 43] show that a single FFN layer in pretrained LLMs can be converted to an MoE layer, by splitting the matrices of FFN layers into exclusive sub-matrices and employing a router to activate only neurons in a subset of experts.
+Similar to MoE models, MoEfication can effectively accelerate LLM inference by only using part of the parameters [43, 20].
+Model Weight Sparsity.
+Another orthogonal direction regarding model sparsity is static model weight sparsity [7, 33], which often prunes the model weights and keeps them static during inference.
+For instance, Wanda [33] estimates model weight importance and prunes unimportant weights, so all inputs will use the same subset of weights.
+In contrast, contextual sparse models select a different subset of weights for different inputs.
+Nonetheless, model pruning can be applied to contextual sparse models too, meaning a complementary optimization to contextual sparsity.
 3 Background and Motivation
 In this section, we first briefly introduce the MoEfication of FFN layers in transformers.
-Then, we present a study on the limitation of applying noisy top-K Softmax routing to train transformers for better sparsity.
+Then, we present a study on the limitation of applying noisy top-K Softmax routing in the Moefication setting.
 3.1 Background: MoEfication
-FFN layers occupy 60% of the total model computation (Liu et al., 2023) but exhibit high activation sparsity.
-In contrast to unstructured sparsity methods like weight pruning, activation sparsity in ReLU-based FFN layers is more hardware-friendly, since intermediate neurons in FFN layers can be grouped as structured chunks (Mirzadeh et al., 2023; Liu et al., 2023).
-MoEfication (Zhang et al., 2022b) is a way to group functional partitions in FFN layers, thereby converting FFN layers to MoE layers for better utilizing the activation sparsity.
-As known, if a transformer’s hidden states and FFN intermediate dimension are and , respectively, the FF layers process the input as following:
-where , are weight metrics, are biased term and is an activation function111SwiGLU in LLaMA needs two matrices to calculate , and the neurons in FFNs can still be used to form experts..
+MoEfication [43] is a way to group neurons in FFN layers, thereby converting FFN layers to MoE layers for better execution speedup.
+For a transformer whose hidden states and FFN intermediate dimension are and , respectively, the FFN layers process the input as follows:
+where , are weight metrics, are biased term and is an activation function.
 MoEfication aims to group neurons in the FFN intermediate layer into experts.
-Then, a router is trained for each FFN layer to select most important experts to improve inference sparsity.
-A recent work, Deja Vu (Liu et al., 2023), also uses a similar setting but treats each neuron as an expert.
+Then, a router is trained for each FFN layer to select experts (k < n), thus reducing activation load, to speed up inference.
+A recent work, Deja Vu [20], uses a similar setting but treats each neuron as an expert.
 However, SOTA MoEfication methods overlook the potential for further optimizing the activation sparsity of LLMs.
-In this paper, besides optimizing LLMs for task performance, we propose to optimize models for fast inference as well (Figure 1).
-By designing efficiency-related loss, we find that our novel training method drives LLMs to achieve better sparsity with marginal performance drops.
-Besides, for LLMs employing soft activations such as GeLU (Hendrycks & Gimpel, 2016) and SwiGLU (Shazeer, 2020), MoEfication (Zhang et al., 2022b) proposes to replace soft activations with ReLU and fine-tune the model to improve activation sparsity.
-Although recent works (Zhang et al., 2022b; Mirzadeh et al., 2023) show that replacing soft activation functions with ReLU only has a marginal impact on model performance, our evaluations indicate that this conclusion is not universally applicable across all datasets.
-It also is crucial to develop methods for effectively utilizing sparsity in LLMs with soft activation functions.
+Besides, for LLMs employing soft activations such as GeLU [11] and SwiGLU [29], MoEfication [43] proposes to replace soft activations with ReLU and fine-tune the model to improve activation sparsity.
+Although recent works [43, 24] show that this replacement has a marginal impact on model performance, we notice that this hypothesis is not widely applicable (Section 5.2).
 3.2 Limitations of Noisy Top-K Softmax Routing
-A potential choice to train the model to adapt to MoE layers is noisy top-K Softmax routing (Fedus et al., 2022; Shazeer et al., 2016), which is a commonly adopted approach to train MoE models from scratch, like Switch Transformers (Fedus et al., 2022).
-This method entails selecting experts based on the highest router outputs and then calculating Softmax values for these selected outputs.
-The output of the MoE layer is determined by summing the products of these softmax values with their corresponding experts’ outputs, which is critical for differentiating router parameters.
+One potential solution to train and convert the model into MoE layers is noisy top-K Softmax routing [6, 31].
+It selects experts based on the highest router outputs and then calculates Softmax values for these selected outputs.
+The output of the MoE layer is determined by summing the products of these softmax values with their corresponding experts’ outputs.
 Additionally, a small Gaussian noise is added to the router’s outputs during training to encourage exploration across different experts, which addresses the issue of unselected experts being non-differentiable.
-However, our empirical findings surprisingly indicate that noisy top-K does not converge well and encounters a biased expert score issue in MoEfication tasks.
-To study the effectiveness of noisy top-K Softmax routing, we follow the standard MoEfication setting (Zhang et al., 2022b) to form experts in FFN layers: we fine-tune a model on a downstream dataset and group neurons into experts via parameter clustering.
-Subsequently, we integrate Softmax routers into models and jointly train both the model and routers with noisy top-K routing.
-The evaluation results on two datasets are shown in Figure 2.
-Surprisingly, we observe similar accuracy drops across all sparsity levels, including at very low levels of sparsity.
-To study the reason behind this, we analyze the expert scores produced by routers and uncover a biased expert issue in MoE layers: one expert in each MoE layer receives an expert score close to 1, while the expert scores for all other routers are nearly 0.
-This means that only one expert contributes to the inference in each MoE layer, which causes the performance drops.
-Before MoEfication, expert outputs in FFN layers can be considered to be multiplied by an expert score equal to 1.
-The sum of expert scores in Softmax routing is constrained to 1, which results in many experts’ outputs being multiplied by expert scores significantly smaller than 1.
-Meanwhile, fine-tuning LLMs involves a relatively low learning rate and fewer training iterations.
-This makes it difficult for neurons in FFN layers to adapt to such largely reduced expert scores.
-Consequently, the model tends to rapidly converge towards allocating the entire Softmax budget to the most important expert, assigning other experts with expert scores close to 0.
-Another limitation of the noisy top-K routing approach is its requirement for a fixed, pre-defined number of selected experts for all inputs and layers.
-However, the sparsity level can vary depending on the specific inputs and layers.
-A more adaptive expert selection strategy helps achieve a better trade-off between model efficiency and quality.
+To study the effectiveness of noisy top-K Softmax routing, we follow the standard MoEfication setting [43] to form experts in FFN layers: we fine-tune a model on a downstream dataset and group neurons into experts via parameter clustering.
+Subsequently, we integrate Softmax routers into models and jointly train both the model and routers.
+As shown in Figure 2, we observe noticeable accuracy drops even at very low sparsity levels.
+When diving into the expert scores produced by routers, we observe a very biased expert score allocation: only one expert in an MoE layer has an expert score close to 1, while other experts receive nearly 0 scores.
+This results in only one expert contributing to the inference in each MoE layer, leading to performance drops.
+A potential reason behind this biased allocation is that the sum of expert scores in Softmax routing is constrained to 1.
+Unlike traditional MoE models, which typically select 1-2 experts [30, 6], MoEfied models need to select a much larger number of experts [20, 43], but the sum of experts score is still 1; this can complicate the allocation of the Softmax budget and cause a biased allocation.
 4 Methodology: Learn To be Efficient
-In this section, we present, LTE (Learn To be Efficient), to learn more structured MoE in FFN layers for fast inference.
-Challenges. Converting a pre-trained LLM’s MLP block into a structured Mixture of Experts (MoE) leads to three practical challenges:
-1.
-How to group neurons into experts?
-Since neurons in the same expert are always activated together, A smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
-2.
-How to train routers more stably?
-The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2).
-How to jointly train routers and model in a MoEfication setting is still under-explored.
-3.
-How to select the right experts in serving?
-The number of experts needed can depend on specific inputs and layers.
-A more adaptive selection strategy leads to a better trade-off between model efficiency and quality.
+In this section, we present LTE (Learn To be Efficient) to learn more structured MoE in FFN layers for fast inference.
+Specifically, we tackle two key challenges toward practical MoEfication for diverse LLMs: (1) How to train routers more stably?
+(Section 4.2); and (2) How to select the right experts in serving?
+(Section 4.2.2).
 4.1 Experts Grouping
+The first step of LTE is to group neurons in FFN layers to form experts.
 Expert grouping aims to group neurons that are often activated together, thereby improving efficiency by reducing the total number of activated neurons.
@@
 Moreover, we need to group neurons without a significant performance drop on the pre-trained model.
-Following the previous work (Zhang et al., 2022b), we use parameter clustering to split MLP neurons into different experts.
+Following the previous work [43], we group 32 neurons as an expert in FFN layers, and use parameter clustering to group MLP neurons into different experts.
 Fundamentally, in FFN layers, each neuron is associated with a column of (which is a vector).
-Parameter clustering split first treats the corresponding vector in as the feature for a neuron, then applies the balanced K-Means algorithm (Malinen & Fränti, 2014) to cluster neurons into clusters.
-Each cluster has neurons by default and will be treated as an individual MoE expert.222 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: We use the weight of the gate network, , to cluster neurons.
+Parameter clustering first treats the corresponding vector in as the feature for a neuron, then applies the balanced K-Means algorithm [22] to cluster neurons into clusters.
+Each cluster has neurons by default and will be treated as an individual MoE expert.111 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: We use the weight of the gate network, , to cluster neurons.
+We also conduct an ablation study on other alternative grouping strategies in Section 5.4.
 4.2 Adaptive Expert Routing
-After allocating neurons to different experts, we need to train and orchestrate routers to select experts for input tokens.
-4.2.1 Router structure
-Similar to existing MoE designs, we use a single fully-connected layer as the router network.
+4.2.1 Router Design
+After constructing the experts, we need to decide on the right routing strategy.
+Similar to existing MoE designs, we use a fully-connected layer as the router network.
 The input to the router is the output of the preceding self-attention block, and the router output is the expert score for each expert.
-Routing function. To address the biased expert score issue (Section 3.2), instead of using conventional Softmax routing, we employ the Sigmoid function on router outputs to get expert scores:
+Routing function.
+To address the biased expert score issue (Section 3.2), we employ the Sigmoid function on router outputs to get expert scores.
+Unlike the Softmax function, the Sigmoid function computes each expert score independently, which circumvents the biased expert score due to the constraint on the sum of outputs in the Softmax function:
 where is the input to the corresponding FFN layer, is the index for the expert, and is the router network weight for the -th expert.
-Unlike the Softmax function, the Sigmoid function computes each expert score independently, which circumvents the biased expert score due to the constraint on the sum of outputs in the Softmax function.
 Threshold-based experts selection.
-SOTA methods (Zhang et al., 2022b; Fedus et al., 2022) select a fixed number of experts for all layers and inputs.
+SOTA methods [43, 6] select a fixed number of experts for all layers and inputs.
 However, the necessary number of experts may differ depending on inputs and layers.
-Here, we propose to select experts more adaptively for different inputs and layers for a better trade-off between model sparsity and quality.
-Specifically, we leverage a threshold-based method to enable adaptive selection:
+Here, we propose to select experts more adaptively for different inputs and layers for a better trade-off between model sparsity and quality, by leveraging a threshold-based method to enable adaptive selection.
+The MoEfied FFN layers become:
 where is the output for the -th expert, is the indicator function, and is a predefined threshold.
-Only the experts whose score is larger than will be activated.
-Threshold-based routing does not have to select a fixed number of experts in each MoE layer.
-Instead, it selects experts with sufficiently large expert scores.
+Only experts with score larger than will be activated.
 Consequently, within the same MoE layer, as the router generates larger expert scores, more experts are selected.
 4.2.2 Two-stage LTE Training
-In practice, training the threshold-based Sigmoid router leads to three challenges.
-Firstly, Sigmoid function considers individual experts independently, making it hard to tease apart important experts from the rest.
-Secondly, the threshold-based indicator function in Equation 2 is non-differentiable, thereby preventing the training on routers.
-Lastly, manually choosing a threshold for each MoE layer can be impractical.
+Training the router poses three practical challenges.
+Firstly, as we consider training experts independently to mitigate bias (e.g., using a Sigmoid routing function), it is difficult to tease apart important experts from the rest.
+Secondly, the non-differentiability of threshold-based expert selection further complicates router training.
+Lastly, setting thresholds for each MoE layer is non-trivial.
 To tackle these challenges, we next propose a novel two-stage training algorithm with an efficiency loss penalty to MoEfy LLMs.
-As shown in Algorithm 1, our training algorithm consists of two training stages: a) model-router training stage and b) model adaption stage.
+Our training algorithm consists of two training stages: a) model-router training stage and b) model adaption stage.
 Stage 1: Model-router training.
-The goal of the model-router joint training stage is to train Sigmoid routers to capture the importance of experts for given inputs.
-In this stage, we jointly train routers and model parameters.
-To address the non-differentiability issue, we switch the expert selection to a “soft” mode in this stage:
+In this stage, we jointly train routers and model parameters to capture the importance of experts for given inputs in the Sigmoid routers.
+To address the non-differentiability issue, we switch the expert selection to a “soft” mode:
 Instead of selecting a subset of experts, we always select all experts and multiply expert outputs with the corresponding expert score to make both router and model differentiable.
-Since there is no discrete selection in MoE layers, all parameters are trained for each input token.
+Since there is no discrete selection in MoE layers, all parameters are trained for each iteration.
 Compared to Softmax routers, Sigmoid routers score each expert independently and do not introduce any competition on expert scores among experts.
@@
 As such, we design an efficiency loss penalty, , to introduce competition among experts for Sigmoid routers:
+where is the number of layers, and is the number of experts in each layer.
 The efficiency loss calculates the mean of expert scores across all layers.
 This efficiency loss penalizes the output magnitudes of routers, driving routers to selectively allocate lower expert scores to less important experts, which helps distinguish experts with different importance.
-Moreover, instead of assigning the same sparsity for all layers, the efficiency loss also induces inter-layer competition, allowing LTE models to allocate adaptive sparsity for different layers.
-Our analysis in Figure 6 shows that LTE models are capable of adaptively allocating sparsity across different layers.
+Moreover, instead of assigning the same sparsity for all layers, the efficiency loss also induces inter-layer orchestration, allowing LTE-trained models to allocate adaptive sparsity for different layers.
+We show that LTE models are capable of adaptively allocating sparsity across different layers (Figure 15).
 For the threshold in Equation 2, instead of choosing a threshold for each router, we propose to select a predetermined fixed threshold and then train models to fit this threshold, i.e., train models to be threshold-aware.
 Specifically, we use a threshold separability regularizer to drive models to make expert scores more separable for this given threshold:
-where is the number of layers, and is a predefined threshold for all routers (we set for all experiments).
+where is a predefined threshold for all routers (we set for all experiments).
 The separability loss encourages router outputs to diverge from the threshold , which makes the router outputs more separable.
 An example is illustrated in Figure 3.
-The only difference between the models in the two figures is if the model is trained with the separability loss.
-We find that, when training without the separability loss, the expert scores are hard to separate, making it hard to decide the threshold.
+The only difference between the models in the two figures is if the model is trained with a separability loss.
+We find that the expert scores are not separable when training without the separability loss, making it hard to decide the threshold.
 However, training with the separability loss markedly improves the separability of expert scores, allowing us to choose experts with the predefined threshold.
-Combined with the task performance loss , our training loss in model-router training state is:
-The efficiency coefficient, , is used to control the trade-off between inference efficiency and task performance. A higher indicates a larger inference cost penalty, leading to a more sparse model.
-Different from the pretrained LLMs, MoE routers are randomly initiated. We use a larger learning rate (0.01) for routers’ parameters to make routers converge faster.
+Combined with the task performance loss , our training loss in model-router training stage is:
+The efficiency coefficient, , is used to control the trade-off between inference efficiency and task performance, and can be used to control the sparsity of trained models.
+A higher indicates a larger inference cost penalty, leading to a more sparse model.
+is the separability loss, we set for our evaluation.
+We conduct an ablation study on both hyperparameters in Section 5.4.
 Stage 2: Model Adaptation.
@@
 In the model adaptation stage, we freeze the router parameters, switch routers to discrete selection mode (Equation 2), and fine-tune the model to adapt to the discrete routing with the task performance loss: .
+4.3 Hardware-efficient Implementation
+As illustrated in Figure 4, same as Deja Vu [20] and moefication [43], LTE provides very structured sparsity.
+After selecting neurons in intermediate layers, a CUDA kernel needs only to load the relevant columns and rows of and from GPU global memory to SRAM to compute dot product, thus reducing memory-I/O as well as computational overheads.
+Thanks to the structured sparsity provided by LTE, the kernel avoids non-coalesced memory access by storing column-major and row-major .
+In this paper, we use Triton 2.3.0 to implement a customized MLP layer to translate the sparsity to wall-clock time latency reduction.
+The evaluation of our custom kernel is presented in Section 5.3.
 5 Experiment
-In this section, we evaluate the effectiveness of LTE with different models on various datasets.
-In Section 5.2, the experiment results on NLU tasks show that LTE models achieve 80 - 95% sparsity in FFN layers without performance drop.
-Similarly, LTE also outperforms all other baselines on NLG tasks.
-For instance, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks.
+In this section, we conduct extensive evaluationS to verify the effectiveness of LTE.
+For instance, in Section 5.2, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks.
+Along with our hardware-aware custom kernel, LTE reduces 25% LLaMA2-7B inference latency at 50% sparsity.
 5.1 Experimental Setting
@@
 We implement LTE on both encoder-based models (RoBERTa, RoBERTa) and decoder-based models (GPT2-Medium, and LLaMA-7B) with the HuggingFace’s transformers.
-Following previous work (Zhang et al., 2022b), we set each expert to contain 32 neurons in FFN layers for all models.
-Datasets. We evaluate LTE performance using 11 datasets across two setups:
-Natural Language Understanding (NLU): we evaluate on eight tasks from the GLUE dataset (Wang et al., 2018) including sentiment classification:SST-2 (Socher et al., 2013), entailment: RTE (Dagan et al., 2005), grammatical classification: CoLA (Warstadt et al., 2018) two natural language reference tasks: MNLI (Williams et al., 2018), QNLI (Rajpurkar et al., 2016), and three tasks measuring sentence similarity: QQP (Iyer et al., 2017), STS-B (Cer et al., 2017), MPRC (Dolan & Brockett, 2005).
+Following previous work [43], we set each expert to contain 32 neurons in FFN layers for all models.
+Datasets: 1) Natural Language Understanding (NLU): we evaluate on eight tasks from the GLUE dataset [37] SST-2 [32], RTE [4], CoLA [39], MNLI [40], QNLI [28], QQP [12], STS-B [3], and MPRC [5].
 We report the Pearson correlation coefficient for STS-B, the Matthews correlation coefficient for CoLA, and the accuracy for the other six datasets.
-Natural Language Generation (NLG): we evaluate on data-to-text (E2E Challenge (Novikova et al., 2017)), text-to-text summarization (XSum (Narayan et al., 2018)), and next token prediction language modeling (Wikitext103 (Merity et al., 2016)).
-For both E2E and XSum, we report the ROUGE-L (Lin, 2004) to measure the generation text quality and report the perplexity performance for the Wikitext dataset.
+2) Natural Language Generation (NLG): we evaluate LTE on E2E [26], XSum [25], and Wikitext103 [23].
+For both E2E and XSum, we report the ROUGE-L [18] to measure the generation text quality and report the perplexity performance for the Wikitext dataset.
+3) Instruction Tuning: besides downstream tasks, we also evaluate LTE on instruction tuning tasks to evaluate LTE’s generalization capabilities.
+We use Tulu dataset [38] to perform instruction tunning with LTE, and we evaluate models with MMLU benchmark [10]).
+MMLU benchmark is a comprehensive evaluation dataset designed to evaluate the generalization capabilities of LLMs and consists of 15908 questions from 57 distinct tasks.
 Baselines.
-(1) MoEfication (Zhang et al., 2022b; Liu et al., 2023): MoEfication proposes to replace soft activations with ReLU and then fine-tune the model on downstream datasets.
+We compare two baselines with LTE in our paper: (1) MoEfication [43]: MoEfication proposes to replace soft activations with ReLU and then fine-tune the model on downstream datasets.
 In our evaluation, we report the number of MoEfication with parameter clustering and ground-truth expert selection.
-Note that ground-truth selection is not practically feasible but can be considered as an upper bound for expert selection.
-(2) Keep-the-Largest-Activation (KLA): To explore the sparsity upper-bound in LLMs trained with task performance loss alone, we manually create an upper-bound baseline based on an ideal scenario.
-We assume that an oracle can always predict the activation value of each neuron with no cost.
-Given a sparsity level, the oracle always selects neurons with the largest absolute activation values.
-Although KLA is not feasible in real-world settings, it represents an upper-bound performance for models that do not optimize efficiency in training.
-The models used in KLA are also models fine-tuned on downstream tasks.
-5.2 Performance Comparison
+(2) Deja Vu [20]: Deja Vu employs predictors to estimate the values of neurons in intermediate layers, pruning neurons with absolute magnitudes (absolute values for non-ReLU activations).
+The models used in Deja Vu evaluation are models fine-tuned on specific tasks as well.
+5.2 End-to-end Performance Comparison
+LTE, MoEfication, and Deja Vu provide the same type of structured sparsity in FFN layers.
+For the convenience of comparison, we use FFN neuron sparsity as a measure of efficiency in this section.
+We will further discuss how our hardware-efficient implementation translates this sparsity into wall-clock time latency reduction in Section 5.3.
+For our method (LTE), we get models with different sparsity by tuning efficiency loss in Equation 6.
 No performance drop with 80-95% sparsity on NLU tasks.
-We first evaluate our methods on eight NLU Tasks in GLUE datasets(Wang et al., 2018) with two encoder-based models (RoBERTa and RoBERTa).
-Due to the space limit, we report four datasets (MRPC, SST2, QNLI, and MNLI) in Figure 4 and other four datasets (RTE, CoLA, STSB, and QQP) in Figure 8 in Appendix B.
-As shown in Figure 4, our method (LTE) consistently outperforms the baseline models on all datasets.
-Both MoEfication and KLA exhibit performance drop when sparsity reaches 70%–80%, while LTE maintains good performance even at higher sparsity levels (90%).
-The major reason behind this performance gap is that, although LLMs naturally have high sparsity, they are still not optimized for efficiency.
-The efficiency loss penalty in LTE training loss effectively enables LLMs to activate fewer neurons while minimally impacting task performance.
+We first evaluate our methods on eight NLU Tasks in GLUE datasets[37] with two encoder-based models (RoBERTa and RoBERTa).
+Due to the space limit, we report four datasets (MRPC, SST2, QNLI, and MNLI) on RoBERTa in Figure 5 and other results in Figure 13 in Appendix C.
+As shown in Figure 5 and Figure 13, our method (LTE) consistently outperforms the baselines on all datasets.
+MoEfication exhibits a performance drop when sparsity reaches 70%–80%, and Deja Vu experiences performance drop at a relatively low sparsity level (e.g., 30%), while LTE maintains good performance even at higher sparsity levels (90%).
 Moreover, we notice that while ReLU-based models achieve high sparsity with a slight decrease in performance, replacing GeLU with ReLU in RoBERTa can affect performance on some datasets.
-For instance, replacing GeLU with ReLU in a RoBERTa model results in a 10% accuracy drop on the MRPC dataset.
+For instance, replacing GeLU with ReLU in a RoBERTa model results in around 10% accuracy drop on the MRPC dataset.
 50% FLOPs saving on NLG tasks.
 We next evaluate LTE with decoder-based models on three types of generation tasks.
-Similar to NLU tasks, Figure 5 shows that our method outperforms all other baselines.
-NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level compared to NLU tasks.
-However, LTE-trained models have better resilience to the sparsity increase.
-Especially at a high sparsity level, other baselines fail to generate meaningful content, but LTE still generates content with a marginal quality drop.
-In NLG tasks, we also observe that replacing soft activation with ReLU can hurt model performance.
-However, training with ReLU-LLaMA largely mitigates some of the performance drops.
-| XSum | E2E | WikiText-103 | |
+Similar to NLU tasks, Figure 6 shows that LTE outperforms all other baselines.
+NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level
+| XSum | E2E | Wiki | |
 | Full | 12.06 | 12.06 | 12.06 |
-| KLA | 7.14 | 6.42 | 7.87 |
+| Deja Vu | 7.92 | 6.42 | 7.92 |
 | MoEfication | 10.45 | N/A | N/A |
@@
 | LTE (Ours) | 5.38 | 4.65 | 6.59 |
-To better understand the computation saving of LTE, we compare the FLOPs per token of different methods on LLaMA, allowing an up to 0.05 ROUGE-L decrease for XSum and E2E, and an up to 0.5 perplexity (PPL) increase for WikiText-103.
+compared to NLU tasks.
+Yet, LTE-trained models have better resilience to the sparsity increase.
+Especially at a high sparsity level, while other baselines fail to generate meaningful content, LTE still generates content with a marginal quality drop.
+Similarly, we also observe the performance drop caused by replacing soft activation with ReLU.
+Additionally, to better understand the computation saving of LTE, we compare the FLOPs per token of different methods on LLaMA-7B, allowing an up to 0.05 ROUGE-L decrease for XSum and E2E, and an up to 0.5 perplexity (PPL) increase for WikiText-103.
 It is noteworthy that the FLOPs for routers are included in our method, constituting approximately 1% of the total FLOPs of FFN layers.
@@
 The evaluation results show that LTE provides a 1.83x - 2.59x FLOPs speed-up and gives the most FLOPs savings among all methods.
-Those evaluation results demonstrate that LTE effectively provides higher sparsity for LLMs even with soft activation functions.
-5.3 Ablation Study & Analysis
-Adaptive Sparsity across Layers.
-Instead of assigning a fixed sparsity to each layer, LTE uses the efficiency loss to introduce competition across different layers to allocate the computation budget to more important layers.
-Here, we conduct a study on the distribution of sparsity across layers in models trained with LTE.
-As shown in Figure 6, we observe that LTE achieves different sparsity levels across different layers.
-For example, the MRPC (RoBERTa) model with 0.46 sparsity has 0.2 sparsity at layer 2, but has more than 0.8 sparsity at layers 10 and 11.
-Another intriguing finding is that RoBERTa and GPT2-M have very different sparsity patterns.
-In RoBERTa, the sparsity of MoE layers significantly increases with depth, contrasting with GPT2-M, where MoE layer sparsity decreases as layer depth increases.
-This difference can stem from differences in the design of these two transformers and tasks.
-Given that MRPC is a sentence classification task, encoder-based transformers like RoBERTa primarily rely on the ”[CLS]” token for classification.
-Consequently, other tokens become less important in deeper layers, allowing sparser FFN layers.
-In contrast, GPT2, a decoder-based transformer designed for language generation, requires all tokens to generate the next token.
-Furthermore, recent research (Geva et al., 2021) on interpretability suggests that deeper FFN layers store more complex semantic information, which may drive deep FFN layers to be less sparse.
+LTE Outperforms SOTA on instruction tuning.
+To further evaluate the effectiveness of LTE, we evaluate LTE as a chatbot and verify model performance in a few-shot learning scenario.
+We fine-tune LTE on the Tulu instruction tunning dataset [38], and then evaluate the few-shot performance on MMLU benchmark [10].
+We use LLaMA2-7B as the base model for LTE training.
+For a fair comparison, we also fine-tune base models with the Tulu dataset for other baselines.
+Specifically, for Deja Vu, we use Tulu-fine-tuned LLaMA2-7B as the base model.
+For Moefication, we use Tulu-fine-tuned ReLULLaMA-7B as the base model to ensure the best possible baseline performance.
+We present the 5-shot MMLU accuracy comparison in Figure 7.
+The evaluation results show that LTE outperforms other baselines across all sparsity levels.
+Yet, LTE shows a weaker sparsity-performance trade-off compared to downstream tasks.
+Our hypothesis is that LTE routers change the structure of the original LLaMA model, which influences general language patterns learned by models in pretraining.
+Instruction tuning alone is not enough to fully recover this ability (considering instruction tuning has much smaller datasets).
+We believe that training with a wider variety of data (like pretraining data) will further improve LTE performance.
+5.3 Translate Sparsity to Wall-clock Time Speedup
+25% wall-clock time latency reduction.
+As discussed in Section 4.3, the sparsity introduced by LTE is highly structured and can be easily translated to latency reduction.
+This section demonstrates that our custom Triton kernel effectively translates the LTE sparsity to wall-clock time speed up.
+We evaluate wall-clock time speed up using LLaMA2-7B on a single 3090Ti.
+We also reported vanilla Pytorch indexed matrix multiplication.
+Due to the requirement for new memory allocations, slicing in Pytorch is very time-consuming.
+This makes indexed matrix multiplication even slower than dense matrix multiplication at most sparsity levels.
+We report the wall-clock time latency comparison in Figure 8 (noting that the router inference latency is already included in LTE latency).
+Compared to a dense FFN block, our custom Triton kernel achieves nearly linear speed-up with respect to sparsity (Left).
+At around 50% sparsity, LTE reduces end-to-end generation latency by approximately 25%.
+5.4 Ablation Study
+Experts grouping algorithms.
+Three different expert grouping algorithms are explored in [43]: random grouping, parameter K-means clustering, and co-activation graph split.
+The co-activation graph split algorithm constructs a graph where each neuron represents a node and the edges denote the frequency of co-activation between neuron pairs.
+This graph is subsequently divided into subgraphs using the graph split algorithm [15], with each subgraph representing a group of neurons.
+To study the effectiveness of these grouping algorithms, we compare their LTE performance by fine-tuning GPT2 on the WikiText-103 dataset.
+As shown in Figure 11, both co-activation graph split and parameter K-means grouping have very similar performance (similar to finding in [43]) and outperform random grouping.
+Given that parameter K-means grouping is simpler to implement, as co-activation graph split requires collecting activation data, we use K-means grouping in our paper.
+The relation between efficiency loss hyperparameter and sparsity.
+As discussed in Section 4.2.2, efficiency loss hyperparameter is used to balance the trade-off between task performance and sparsity (efficiency).
+We illustrate the relation between and sparsity across different models and tasks in Figure 11.
+For a given task and model, increasing results in greater sparsity.
+Ablation study on separability loss.
+Another hyperparameter in LTE training, , is used to encourage the separability of router outputs, which facilitates us to choose the threshold to pick neurons (as shown in Figure 3).
+We compare models trained with different lambda in Figure 11.
+When increasing from (i.e., no separability loss) to , we observe a constant perplexity drop for all sparsity.
+However, if we keep increasing , perplexity does not further decrease.
 Effectiveness of LTE-trained routers.
@@
 We believe that our work can inspire more research on designing more efficiency-aware training methods, making LLMs more accessible to the broad community.
-Impact Statement
+Broader Impact
 This paper presents the Learn-To-be-Efficient (LTE) algorithm to accelerate the LLM inference.
@@
 To sum up, the LTE algorithm not only advances the field technically but also broadens the scope of AI’s benefits to a wider, more inclusive community.
-Appendix A Experiment Settings
+Limitations
+While LTE outperforms other baselines, LTE needs further fine-tuning to get contextual sparse models.
+(Yet, training is a one-time investment that yields long-term benefits for LLM serving.) A future study is to combine LTE with other parameter-efficient fine-tuning techniques to reduce LTE training overhead.
+Another limitation is our computational resource, which prevents us from training LTE on pretraining or RLHF.
+As discussed in Section 5.2, we believe that training with a wider variety of data will further improve LTE performance.
+We leave this for future exploration.
+Appendix A Code and Datasets License
+Codebase. Our model implementation is based on Huggingface transformer repo: Apache-2.0 license
+Datasets. We list the license of used datasets as follow:
+GLUE dataset [37]: Custom License;
+WikiText103 [23]: CC BY-SA 3.0;
+XSum [25]: MIT License;
+E2E [26]: CC4.0-BY-SA;
+MMLU [10]: Custom License.
+Appendix B Experiment Settings
 We presented all training hyperparameters in Table 2, 3, 4, and 5.
 For hyperparameters presented in {}, we select the best hyperparameter for each task.
-We follow the settings in RoBERTa paper (Liu et al., 2019) to fine-tune RoBERTa on GLUE datasets.
+We follow the settings in RoBERTa paper [19] to fine-tune RoBERTa on GLUE datasets.
 We set the coefficient for the separability loss ( in Equation 6) to be 0.5 for all stage 1 training.
-Appendix B Additional Evaluation Results
+We use different to control the sparsity of trained models (Figure 11).
+Hardware: All models are trained and evaluated on A100, A40, and 3090Ti, depending on memory usage and availability.
+Appendix C Additional Evaluation Results
+C.1 NLU Performance Comparison
 In Figure 13, we present the evaluation results of the rest four NLU datasets.
@@
 The size of CoLA is relatively small and may not be efficient for models to learn good routers.
+C.2 Further Analysis on Sparsity Provided by LTE
+Contextual Sparsity for Larger Batches.
+Similar to previous contextual sparsity work [20], we observe that as the batch size increases, more neurons are activated, leading to a reduction in union sparsity.
+These dynamics are illustrated in Figure14.
+Similar to Deja Vu [20], the number of activated neurons does not grow linearly with the batch size, indicating a non-uniform distribution of parameter access from different inputs.
+This property can extend LTE for a high-throughput batch setting by batching inputs activating similar neurons together to achieve a high union sparsity.
+Adaptive Sparsity across Layers.
+Instead of assigning a fixed sparsity to each layer, LTE uses the efficiency loss to introduce competition across different layers to allocate the computation budget to more important layers.
+Here, we conduct a study on the distribution of sparsity across layers in models trained with LTE.
+As shown in Figure 15, we observe that LTE achieves different sparsity levels across different layers.
+For example, the MRPC (RoBERTa) model with 0.46 sparsity has 0.2 sparsity at layer 2, but has more than 0.8 sparsity at layers 10 and 11.
+Another intriguing finding is that RoBERTa and GPT2-M have very different sparsity patterns.
+In RoBERTa, the sparsity of MoE layers significantly increases with depth, contrasting with GPT2-M, where MoE layer sparsity decreases as layer depth increases.
+This difference can stem from differences in the design of these two transformers and tasks.
+Given that MRPC is a sentence classification task, encoder-based transformers like RoBERTa primarily rely on the "[CLS]" token for classification.
+Consequently, other tokens become less important in deeper layers, allowing sparser FFN layers.
+In contrast, GPT2, a decoder-based transformer designed for language generation, requires all tokens to generate the next token.
+Furthermore, recent research [9] on interpretability suggests that deeper FFN layers store more complex semantic information, which may drive deep FFN layers to be less sparse.
 | Hyperparameter | LTE-Stage 1 | LTE-Stage 2 |
@@
 | Warm up ratio | 0.06 | 0.06 | 0.06 |
+| Hyperparameter | Fine-tuning | LTE-Stage 1 | LTE-Stage 2 |
+| Learning rate | 2e-5 | 2e-5 | 2e-5 |
+| Training batch size | 128 | 128 | 128 |
+| Training epochs | 3 | 1 | 4 |
+| Weight decay | 0 | 0 | 0 |
+| Warm up ratio | 0.06 | 0.06 | 0.06 |